OpenStudy (deoxna):

The following is a derivative, what would be an equation modeling it and the original function (integral)?

OpenStudy (deoxna):



OpenStudy (anonymous):

1/(x^2+1)

OpenStudy (anonymous):

is the max on the y-axis or in first quadrant as shown by your graph?

OpenStudy (deoxna):

In the first quadrant

OpenStudy (anonymous):

ok so just do a horizontal translation 1 unit to the right... 1/[(x-1)^2+1]

OpenStudy (deoxna):

Ok, but how would you integrate that?

OpenStudy (anonymous):

that's the arctan formula... \[\int\limits_{?}^{?}\frac{dx}{a^2+x^2}=\frac{1}{a}\arctan(\frac{x}{a})+C\]

OpenStudy (deoxna):

Wierd, I had never heard of that.

OpenStudy (anonymous):

really?, you've never taken the derivative of arctan x? derivative of arctan x is 1/(x^2 + 1)

OpenStudy (deoxna):

Nope, but thanks anyway!